# Spectrum of slip dynamics, scaling & statistical laws emerge from simplified model of fault and damage zone architecture

### v1

@@ 2 Fault Volume Model
This study focuses primarily on the impact of the geometry and architecture of fault zone on slip dynamics. Hence, this fault volume model is a simplified, canonical, representation of a fault and its damage zone. We begin by defining the elastic properties of the medium, its loading and initial conditions. All length scales are non-dimensionalized using frictional length scales (either the nucleation length or the cohesive zone size) computed using the initial stress state of the medium. These length scales are clearly approximations, as they evolve during the simulation as the the normal traction evolves due to loading and interaction. We then establish the frictional properties of the main fault and set its length to four times the nucleation length. Next, we define a damage zone with a width five times the cohesive zone size of the main fault. Within this damage zone, we assume a hierarchical distribution of off-fault fractures. The density of these fractures decreases according to a power law with fixed exponent and maximum value. The size distribution also follows a power law with a fixed exponent, where the smallest length scale is approximately 1m and the largest is on the order of the main fault length. We select the orientation distribution of these fractures from one of four possible configurations. We assume each off-fault fracture is twice its nucleation length to determine the critical slip distance, $d_{c}$, for each fracture. This ensures $d_{c}$ scales automatically with fault length as described above. Finally, we statistically sample five times from the distributions (length, orientation, and density) to generate five samples of the fault volume model for a given orientation of the off-fault fractures and $a/b$ ratio. We thus generate a total of 40 different fault volume models for this study each of them including around 700-1000 faults. In the following sections we describe each of the steps in more detail.
@@ 2.1 Geometry of the main fault and its damage zone
Figure 2: Fault volume geometry of a case study: a) sketch of fault volume geometry (not to scale): the main fault in black and the off-fault fractures in red b) off-fault fracture density perpendicular to the main fault (linear scale) c) probability density function of length distribution of the faults, including the main fault d) rose diagram of off-fault fracture orientations and principal stress direction e) critical slip distance of all fractures with respect to their length
The fault volume in the model comprises a main fault exhibiting right-lateral movement in an isotropic linear elastic medium ($\mu=30$ GPa, $\nu=0.25$), with fixed frictional properties, accompanied by off-fault fractures within a damage zone surrounding the main fault. Since geometric roughness is a characteristic observed at multiple scales in natural faults [149, 172, 105, 152, 34, 35], in this fault volume model, the main fault is characterized by self-similar roughness with $\alpha=0.002$ [61]. Its extent is four times its nucleation length $L_{\text{main}}=4L_{\text{nuc}}^{\text{main}}$ (Figure 2a). Note that the nucleation length is just an estimation, calculated for a single planar infinite fault, following [165, 199] (See A8 for more details). We set the damage zone width as $W=5L_{b}^{\text{main}}$ making it proportional to the main fault’s cohesive zone size ($L_{b}^{\text{main}}$). This choice reflects the fact that the cohesive zone size governs the decay rate of stress perturbations from the main fault, which in turn controls the characteristic length scale over which main and off-fault fractures interact [137]. We also add a fracture damage zone at each end of the main fault whose fracture density is assumed to be uniformly distributed.
@@
We next set the orientation statistics of the off-fault fractures. We experiment with four different orientations of fractures. In the first three cases, the fractures are deliberately well oriented to maximize reactivation and thus interaction with the main fault. We examine favorable orientations for right-lateral failure, left-lateral failure, and a combination of both (see example in Figure 2d). Optimal planes make an angle of $\theta=\pm\left({\pi}/{4}-{1}/{2}\tan^{-1}f_{0}\right)$ with the direction of maximum principal stress [23], where $f_{0}$ is the reference friction coefficient. Here $f_{0}=0.6$, so $\theta\approx\pm 30^{\circ}$. The initial maximum principal stress makes an angle of 120with the horizontal axis. In this study, we will consider off-fault orientation distributions in the ranges [90± 10] and [150± 10], respectively. For the final case, the fractures run parallel to the main fault, as it has been observed that the growth process of faults reveals that off-fault cracks tend to be oblique by [0–20] to the trace of major faults [147]. For this case, we consider the fracture orientations in the range [0± 10].
Observations from fault zones consistently show a significant decrease in the density of off-fault cracks with increasing distance from the main fault. This observation has been well documented in fault zones [44, 65, 150, 157]. Based on these observations of damage zones surrounding major faults, we model a power-law decay of fracture density with distance normal to the main fault, using an exponent of 1 (Figure 2b). We fix the peak density of off-fault fractures to be around 4 fractures per square meter, which is consistent with observations [44, 124, 66, 157]. We further assume that all of the off-fault fractures are pure mode II fractures. The main numerical bottleneck in simulating tensile off-fault fractures in an earthquake cycle model is the lack of a good cohesive law that allows for the healing of the cohesive strength after rupture. We therefore restrict ourselves to shear fractures that follow rate-and-state friction law.
@@ 2.2 Frictional Parameters
Friction is assumed to follow regularized rate-and-state friction, spatially uniform and rate-weakening on all faults with $f_{0}=0.6$ (See A2). Specifically, for the main fault, the ratio $a/b$ is fixed at 0.75 and $d_{c}$ is set at 2 mm. Here, $a$ is the direct effect parameter that governs the instantaneous change in friction with a change in slip rate, $b$ is the the evolution effect parameter which controls how friction evolves over time via changes in the state variable slip and $d_{c}$ is the critical slip distance. Since $a/b$ is a crucial parameter in controlling the frictional length scales (see A8) we parametrize the problem by varying this parameter. All off-fault fractures within the model share the same $a/b$ ratio, with tested values of 0.4 and 0.5. The characteristic slip distance ($d_{c}$) is scaled with the length of off-fault fractures (Figure 2e) such that each off-fault fracture is twice the nucleation length.
Earthquake sources are conventionally modeled as shear ruptures occurring on preexisting faults within the seismogenic zone. Nevertheless, these faults are characterized by complex geometric irregularities and mechanical heterogeneities that manifest across a wide range of spatial scales throughout the fault zone. Be it in the lab scale or the field scale, the net weakening distance is a homogenized manifestation of weakening processes occurring at smaller scales [136, 69]. This approach is also inspired based on the observation that the so called fracture energy ($G$’) scales with slip [135, 2], and slip is itself related to fracture length, through elasticity. Furthermore, [165] showed that fracture energy is approximately proportional to $d_{c}$ within the rate-and-state friction framework. These observations together justify scaling $d_{c}$ with fracture length in our model. This also has the additional advantage of keeping the computational cost reasonable. To keep the values of $d_{c}$ realistic, we ensure that the smallest fracture is around 1 m, the length scale of laboratory experiments and use laboratory inferred values of $d_{c}$ [114, 175, 106] of the order of several tens of microns.
@@ 2.3 Loading and Initial conditions
The medium has an initial prestress state with $\sigma_{11}^{0}=-6.46\,\text{MPa}$, $\sigma_{12}^{0}=3.88\,\text{MPa}$, $\sigma_{22}^{0}=-10.94\,\text{MPa}$. It is also loaded with uniform far-field stressing rate, $\dot{\sigma}_{11}^{\infty}=-0.0064\,\text{Pa s}^{-1}$, $\dot{\sigma}_{12}^{\infty}=0.0038\,\text{Pa s}^{-1}$, $\dot{\sigma}_{22}^{\infty}=-0.0108\,\text{Pa s}^{-1}$. This corresponds to strain rates $\dot{\varepsilon}^{\infty}_{11}=-4.93\times 10^{-14}\text{ s}^{-1},\dot{\varepsilon}^{\infty}_{22}=-1.23\times 10^{-13}\text{ s}^{-1}$ and $\dot{\varepsilon}^{\infty}_{12}=6.33\times 10^{-14}\text{ s}^{-1}$. The direction of principal stress for far-field stress rate and initial stress are the same, and is assumed here to make an angle of 120with the horizontal axis. At each fault, $\tau_{0}=\sigma_{ij}^{0}n_{j}s_{i}$ and $\sigma_{0}=\sigma_{ij}^{0}n_{j}n_{i}$, where $\tau_{0}$ and $\sigma_{0}$ are initial shear and normal stress on the fault, $n$ and $s$ are the unit normal and unit tangent vectors respectively. All faults are assumed to be initially at steady state, and initial slip rates considered constant: $V_{\text{init}}=V_{0}=10^{-9}$ m/s where $V_{0}$ is the reference slip rate for rate-and-state friction. The system initiates through the activation of a high slip rate patch located on the main fault. We neglect the first artificial earthquake on the main fault forced by loading, and only study the subsequent cycles after the system loses the memory of the initial conditions. To avoid unusual buildup of normal traction on the fault, we simply cap the normal stress at 10 MPa in an elastic-rigidly plastic sense. This needs to be examined in more detail in future work as it is beyond the scope of this paper. The slip rate is capped at a minimum value of $10^{-20}$ m/s at the end of each time step to avoid unnecessary numerical artifacts as suggested in the SEAS benchmark [88].
@@ 2.4 Numerical Implementation
We implement the fault volume model using a quasi-dynamic boundary integral method with spatial convolution accelerated using hierarchical matrices. We account for the full elastic interaction between all faults in the system, including the main fault and all off-fault fractures, To ensure numerical accuracy, we discretize each of the faults such that at least 5 grid points are within the cohesive zone of each fault. This results in about 30000-50000 unknowns in the system per simulation. Each of the simulations is run until we produce between 8 and 10 major earthquakes on the main fault, which corresponds to between 2 and 4 million adaptive time steps. More details on the numerical implementation can be found in Appendix A.
@@ 2.5 Event Detection and Catalog Generation
The detection and classification of events in the system involve several steps. Events are identified when the maximum slip rate exceeds predefined thresholds, with categorization into “fast” or “slow” depending on the specific threshold exceeded: $10^{-3}$ m/s for fast and $10^{-6}$ m/s for slow. Additional values for the slow rupture threshold ($10^{-7}$, $10^{-8}$, and $10^{-9}$ m/s) are also tested. The effect of this threshold on the catalog and the scaling laws will be discussed in the following sections. Subsequently, events that fall within this threshold range are delineated spatiotemporally through the application of the connected-component labeling (CCL) algorithm, commonly utilized in image processing. The detailed event detection algorithm is presented in Appendix B. This process yields information on rupture length $L_{\text{rup}}$ and duration $T$, which allows us to estimate an average rupture velocity $v_{r}=L_{\text{rup}}/T$. We also estimate for each event the final slip $\delta$ and subsequently the seismic moment as follows: $M_{0}=\mu S\delta$, where $\mu$ is the shear modulus and $S$ is the rupture area. Since the model is 2D, we use an equivalent rupture surface $S$ defined as $\pi L_{\text{rup}}^{2}/4$, assuming a circular rupture. We then estimate the moment magnitude $M_{w}=0.67\log_{10}M_{0}-6.06$ [75]. Finally, we apply a filtering process to the detected events, eliminating those with fewer than 5 time steps and spanning less than 5 grid points.
Figure 3: Time series of the moment rate $\dot{M}_{0}$. The black curve represents the contribution coming from the main rough fault, and the red curve represents the summation of the moment rate released by all the fractures. Panel a) The entire seismic cycle. Top x-axis is absolute time in years, and bottom x-axis is time normalized by recurrence time of earthquakes when considering only the main rough fault, without a damage zone ($\sim 62$ years). Panel b), c) and d) show time snapshots of the seismic cycle to highlight certain events. Here x-axis is relative time for the corresponding time snapshot. The spatiotemporal description of these events are shown in Figure 4.
@@ 3.1 Seismic Cycles and Slip Dynamics
In this section, we provide a concise summary of results based on a case study. The geometric details are visually presented in Figure 2.
Figure 3 depicts the time series of the seismic moment rate released ($\dot{M}_{0}$), with the black curve illustrating the contribution originating from the main rough fault. In contrast, the red curve represents the cumulative moment rate released by all the off-fault fractures combined. In Figure 3a, in order to measure the impact of the fault volume on slip dynamics, we normalize time by recurrence time of earthquakes when considering only a single rough fault, without a damage zone. Notably, the presence of the fault volume introduces a higher frequency of ruptures on the main fault, revealing a significant increase in seismic activity. Specifically, for the fault volume scenario presented here, we observe eight ruptures, twice as many as in the single-fault case, which exhibits only four over the same time period (Figure 3a). This substantial difference emphasizes the significant perturbation introduced by the fault volume on the seismic behavior of the system, resulting in a discernibly shorter recurrence interval. Importantly, the recurrence time is not constant but rather exhibits variability, giving rise to intricate seismic events. On the other hand, while the main fault experiences seismic cycles, we can observe that the off-fault fractures seem to be consistently releasing moment throughout the seismic cycle.
Figure 4: Sequences of slip rate profiles on the main fault. Top x-axis is average slip accumulated during the specific time sequence. Left y-axis represents the along-strike distance on the main fault. Horizontal lines show the specific durations for different phases of the rupture sequence, with blue lines indicating slow phases and red lines indicating fast phases. Nucleation and afterslip phases are indicated by two blue lines at the beginning and end of fast events. The colormap represents slip rate. Black circles represent events detected on the fractures, and projected onto the main fault strike. Yellow stars represent the epicenters of the different events. Panel a) shows complex partial ruptures on the main fault. Panel b) show one slow slip event. Panel c) shows a full rupture on the main fault. Panel d) shows an event, nucleated from one end of the fault and accelerated in a cascading process.
@@
Having explored the temporal slip dynamics, our focus will now shift to examining the spatial complexity of the seismic events recorded on the main fault (Figure 4). In the presented figure, we delve into the spatial complexity of seismic events recorded on the main fault by looking at the slip rate profiles along strike as a function of spatially averaged cumulative slip. In addition, we project all off-fault activity during this event onto the main fault and denote them by simple black circles. For our analysis, we carefully selected several representative events from the case study, encompassing a diverse range of slip occurrences. This includes partial ruptures, full ruptures, and slow slip events, each characterized by distinct magnitudes and durations. For the fast ruptures, we present both the nucleation phase, where the slip rate starts to accelerate from an arbitrarily small threshold, i.e. $10^{-9}$ m/s, prior to reaching some geodetically detectable threshold, i.e., $10^{-6}$ m/s, and the afterslip phase, where the slip rate decelerates below the same limit. Figure 4a features two partial ruptures, each encompassing around half of the main fault. Notably, the second rupture (moment magnitude $M_{w}=4.70$) nucleates at the edge of the first one ($M_{w}=4.56$), demonstrating a cascading effect. Despite the rapid rupture of coseismic phases, which are 17 and 22 seconds, respectively, the nucleation phase lasting 122 days and the afterslip phase taking place over 30 days are significantly slower.

### v2

@@ 2 Fault Volume Model
2.1 Overview of the Fault Volume Model
This study focuses primarily on the impact of the geometry and architecture of a fault zone on slip dynamics. Hence, this fault volume model is a simplified representation of a fault and its damage zone. We begin by defining the elastic properties of the medium, its loading and initial conditions. All length scales are non-dimensionalized using frictional length scales (either the nucleation length or the cohesive zone size) computed using the initial stress state of the medium. We remark that these frictional length scales (see Supporting Information for more details) are merely approximations, as they are based on assumptions of a single planar fault undergoing no change in normal and shear traction over time and always at steady state. We then establish the frictional properties of the main fault and set its length to four times the nucleation length. Next, we define a damage zone with a width, $W$, five times the cohesive zone size of the main fault. Within this damage zone, we assume a hierarchical distribution of off-fault fractures. The density of these fractures decreases according to a power law with a fixed exponent and maximum value. The size distribution also follows a power law with a fixed exponent, where the smallest length scale is approximately 1m and the largest is on the order of the main fault length. We select the orientation distribution of these fractures from one of four possible orientations with respect to the main fault and principal stress directions (Section 2.2). We assume each off-fault fracture is twice its nucleation length to determine the critical slip distance, $d_{c}$, for each fracture. This ensures $d_{c}$ scales automatically with fault length as described above. Finally, we statistically sample five times from the distributions (length, orientation, and density) to generate five samples of the fault volume model for a given orientation of the off-fault fractures and its rate and state friction properties. We thus generate a total of 40 different fault volume models for this study each of them including around 700-1000 faults. In the following sections we describe each of the steps in more detail.
@@ 2.2 Geometry of the Main Fault and its Damage Zone
Figure 1: Fault volume geometry of a case study: a) sketch of fault volume geometry (not to scale): the main fault in black and the off-fault fractures in red b) off-fault fracture density perpendicular to the main fault (linear scale) c) Frequency of length distribution of the faults, including the main fault d) rose diagram of off-fault fracture orientations and principal stress direction e) critical slip distance of all fractures with respect to their length. RL and $L$ stands for right and left lateral slip respectively.
The fault volume in the model comprises a main fault exhibiting right-lateral movement in an isotropic linear elastic medium ($\mu=30$ GPa, $\nu=0.25$), with fixed frictional properties, accompanied by off-fault fractures within a damage zone surrounding the main fault. Since geometric roughness is a characteristic observed at multiple scales in natural faults [139, 161, 98, 142, 30, 31], in this fault volume model, the main fault is characterized by self-similar roughness with $\alpha=0.002$ [53]. Its extent is four times its nucleation length $L_{\text{main}}=4L_{\text{nuc}}^{\text{main}}$ (Figure 1a). Note that the nucleation length is just an estimation, calculated for a single planar infinite fault, following [155, 187] (See Supporting Information for more details). We set the damage zone width as $W=5L_{b}^{\text{main}}$ making it proportional to the main fault’s cohesive zone size ($L_{b}^{\text{main}}$). This choice reflects the fact that the cohesive zone size governs the decay rate of stress perturbations from the main fault, which in turn controls the characteristic length scale over which main and off-fault fractures interact [128]. We also add a fracture damage zone at each end of the main fault whose fracture density is assumed to be uniformly distributed.
@@
We next set the orientation statistics of the off-fault fractures. We experiment with four different orientations of fractures. In the first three cases, the fractures are deliberately well oriented to maximize reactivation and thus interaction with the main fault. We examine favorable orientations for right-lateral failure, left-lateral failure, and a combination of both (see example in Figure 1d). Optimal planes make an angle of $\theta=\pm\left({\pi}/{4}-{1}/{2}\tan^{-1}f_{0}\right)$ with the direction of maximum principal stress [20], where $f_{0}$ is the reference friction coefficient. Here $f_{0}=0.6$, so $\theta\approx\pm 30^{\circ}$. The initial maximum principal stress makes an angle of 120 with the horizontal axis. We thus consider right-lateral fractures (oriented at $\theta=150^{\circ}\pm 10^{\circ}$), left-lateral fractures (oriented at $\theta=90^{\circ}\pm 10^{\circ}$), both right and left-lateral fractures (conjugate planes at $\theta=[90^{\circ};150^{\circ}]\pm 10^{\circ}$). For the final case, the fractures run parallel to the main fault, as it has been observed that the growth process of faults reveals that off-fault cracks tend to be oriented at angles of [0–20] to the trace of major faults [138]. For this case, we consider the fracture orientations in the range [0± 10] (Figure 1d).
Observations from fault zones consistently show a significant decrease in the density of off-fault cracks with increasing distance from the main fault. This observation has been well documented in fault zones [40, 59, 140, 147]. Based on these observations of damage zones surrounding major faults, we model a power-law decay of fracture density with distance normal to the main fault, using an exponent of 1 (Figure 1b shows the sampled distribution). This exponent is higher than values reported in some studies [131, 160] but lower than the values suggested by [140]. We fix the peak density of off-fault fractures to be around 4 fractures per square meter, which is consistent with observations [40, 117, 58, 147]. We further assume that all of the off-fault fractures are pure mode II fractures. The main numerical bottleneck in simulating tensile off-fault fractures in an earthquake cycle model is the lack of a good cohesive law that allows for the healing of the cohesive strength after rupture. We therefore restrict ourselves to shear fractures that follow rate-and-state friction law.
@@ 2.3 Frictional Parameters
Friction is assumed to follow regularized rate-and-state friction, spatially uniform and rate-weakening on all faults with $f_{0}=0.6$ (See Supporting Information for more details). Specifically, for the main fault, the ratio $a/b$ is fixed at 0.75 and $d_{c}$ is set at 2 mm. Here, $a$ is the direct effect parameter that governs the instantaneous change in friction with a change in slip rate, $b$ is the evolution effect parameter which controls how friction evolves over time via changes in the state variable and $d_{c}$ is the critical slip distance. Since $a/b$ is a crucial parameter in controlling the frictional length scales (see Supporting Information for more details) we parametrize the problem by varying this parameter. All off-fault fractures within the model share the same $a/b$ ratio, with tested values of 0.4 and 0.5. The characteristic slip distance ($d_{c}$) is scaled with the length of off-fault fractures (Figure 1e) such that each off-fault fracture is twice the nucleation length.
Earthquake sources are conventionally modeled as shear ruptures occurring on preexisting faults within the seismogenic zone. Nevertheless, these faults are characterized by complex geometric irregularities and mechanical heterogeneities that manifest across a wide range of spatial scales throughout the fault zone. Be it in the lab scale or the field scale, the net weakening distance is a homogenized manifestation of weakening processes occurring at smaller scales [126, 61]. This approach is also inspired by the observation that the so called fracture energy ($G^{\prime}$) scales with slip [127, 2], and slip is itself related to fracture length, through elasticity. Furthermore, [155] showed that fracture energy is approximately proportional to $d_{c}$ within the rate-and-state friction framework. These observations together justify scaling $d_{c}$ with fracture length in our model. This also has the additional advantage of keeping the computational cost reasonable. To keep the values of $d_{c}$ realistic, we ensure that the smallest fracture is around 1m, the length scale of laboratory experiments and use laboratory inferred values of $d_{c}$ [107, 164, 99] of the order of several tens of microns.
@@ 2.4 Loading and Initial Conditions
The medium has an initial prestress state with $\sigma_{11}^{0}=-6.46\,\text{MPa}$, $\sigma_{12}^{0}=3.88\,\text{MPa}$, $\sigma_{22}^{0}=-10.94\,\text{MPa}$. It is also loaded with uniform far-field stressing rate, $\dot{\sigma}_{11}^{\infty}=-0.0064\,\text{Pa\,s}^{-1}$, $\dot{\sigma}_{12}^{\infty}=0.0038\,\text{Pa\,s}^{-1}$, $\dot{\sigma}_{22}^{\infty}=-0.0108\,\text{Pa\,s}^{-1}$. This corresponds to strain rates $\dot{\varepsilon}^{\infty}_{11}=-4.93\times 10^{-14}\text{ s}^{-1},\dot{\varepsilon}^{\infty}_{22}=-1.23\times 10^{-13}\text{ s}^{-1}$ and $\dot{\varepsilon}^{\infty}_{12}=6.33\times 10^{-14}\text{ s}^{-1}$. The direction of principal stress for far-field stress rate and initial stress are the same, and is assumed here to make an angle of 120 with the horizontal axis. At each fault, $\tau_{0}=\sigma_{ij}^{0}n_{j}s_{i}$ and $\sigma_{0}=\sigma_{ij}^{0}n_{j}n_{i}$, where $\tau_{0}$ and $\sigma_{0}$ are initial shear and normal stress on the fault, $\mathbf{n}$ and $\mathbf{s}$ are the unit normal and unit tangent vectors respectively. All faults are assumed to be initially at steady state, and initial slip rates are considered constant: $V_{\text{init}}=V_{0}=10^{-9}$ m/s where $V_{0}$ is the reference slip rate for rate-and-state friction. The system initiates through the activation of a high slip rate patch located on the main fault. We neglect the first artificial earthquake on the main fault forced by loading, and only study the subsequent cycles after the system loses the memory of the initial conditions. To avoid an unusual buildup of normal traction on the fault, we simply cap the normal traction at 10 MPa in an elastic-rigidly plastic sense. This needs to be examined in more detail in future work as it is beyond the scope of this paper. The slip rate is capped at a minimum value of $10^{-20}$ m/s at the end of each time step to avoid unnecessary numerical artifacts as suggested in the SEAS benchmark [81].
@@ 2.5 Numerical Implementation
We implement the fault volume model using a quasi-dynamic boundary integral method with spatial convolution accelerated using hierarchical matrices. To assess the sensitivity of our fault volume model to various parameters, we conducted multiple simulations exploring different fracture orientations and frictional parameters of off-fault fractures. To ensure robustness of our results, each off-fault orientation was subjected to five different statistical samplings, guaranteeing a comprehensive understanding of the system’s behavior across various geometries. Furthermore, we repeated the entire analysis, initially with an $a/b$ ratio of 0.5 for the off-fault fractures, and subsequently with a new $a/b$ ratio of 0.4, while maintaining homogeneous, rate-weakening friction conditions. In total, we tested 40 different configurations with various combinations of off-fault fracture orientations and frictional parameters.
We account for the full elastic interaction between all faults in the system, including the main fault and all off-fault fractures. To ensure numerical accuracy, we discretize each of the faults such that at least 5 grid points are within the cohesive zone of each fault. This results in about 30000–50000 unknowns in the system per simulation. Each of the simulations is run until we produce between 8 and 10 major earthquakes on the main fault, which corresponds to between 2 and 4 million adaptive time steps. More details on the numerical implementation can be found in Supporting Information.
@@ 2.6 Event Detection and Catalog Generation
The detection and classification of events in the system involve several steps. Events are identified when the maximum slip rate exceeds predefined thresholds, with categorization into “fast” or “slow” depending on the specific threshold exceeded: $10^{-3}$ m/s for fast and $10^{-6}$ m/s for slow. Additional values for the slow rupture threshold ($10^{-7}$, $10^{-8}$, and $10^{-9}$ m/s) are also tested. The effect of this threshold on the catalog and the scaling laws will be discussed in the following sections. Subsequently, events that fall within this threshold range are delineated spatiotemporally through the application of the connected-component labeling (CCL) algorithm, commonly utilized in image processing. The detailed event detection algorithm is presented in Supporting Information. This process yields information on rupture length $L_{\text{rup}}$ and duration $T$, which allows us to estimate an average rupture velocity $v_{r}=L_{\text{rup}}/T$. We also estimate for each event the final slip $\delta$ and subsequently the seismic moment as follows: $M_{0}=\mu S\delta$, where $\mu$ is the shear modulus and $S$ is the rupture area. Since the model is 2D, we use an equivalent rupture surface $S$ defined as $\pi L_{\text{rup}}^{2}/4$, assuming a circular rupture. We then estimate the moment magnitude $M_{w}=0.67\log_{10}M_{0}-6.06$ [66]. Finally, we apply a filtering process to the detected events, eliminating those with fewer than 5 time steps and spanning less than 5 grid points so that the catalog properties can be accurately computed. An example of the generated catalog is shown in Figure 2e.
Figure 2: Time series of the moment rate $\dot{M}_{0}$. The black curve represents the contribution coming from the main rough fault, and the red curve represents the summation of the moment rate released by all the fractures. Panel a) The entire seismic cycle. Top x-axis is absolute time in years, and bottom x-axis is time normalized by recurrence time of earthquakes when considering only the main rough fault, without a damage zone ($\sim 62$ years). Panel b), c) and d) show time snapshots of the seismic cycle to highlight certain events. Here x-axis is relative time for the corresponding time snapshot. The spatiotemporal description of these events are shown in Figure 3. Panel e) shows the catalog built from the above continuous moment-rate results using slip velocity threholds as described in section 2.6.
@@ 3.1 Seismic Cycles and Slip Dynamics
In this section, we provide a concise summary of results based on a case study using the fault geometry shown in Figure 1.
Figure 2a depicts the time series of the moment-release rates ($\dot{M}_{0}$), with the black curve illustrating the contribution originating from the main rough fault. In contrast, the red curve represents the sum of the moment-release rates of off-fault fractures. In Figure 2a, in order to measure the impact of the fault volume on slip dynamics, we normalize time by the recurrence time of earthquakes when considering only a single rough fault, without a damage zone. Notably, for the fault volume scenario presented here, we observe eight ruptures, twice as many as in the single-fault case, over the same time period (Figure 2a). This substantial difference emphasizes the significant perturbation introduced by the fault volume on the seismic behavior of the system, resulting in a discernibly shorter recurrence interval. Importantly, the recurrence time is not constant but rather exhibits variability, giving rise to intricate seismic and aseismic events. While the main fault experiences seismic cycles, individual off-fault fractures also undergo their own seismic cycles, but their collective moment release appears relatively steady throughout the main fault’s seismic cycle. We will examine this in more detail in later sections.
@@
Figure 3: Sequences of slip rate profiles on the main fault. Top x-axis is average slip accumulated during the specific time sequence. Left y-axis represents the along-strike distance on the main fault. Horizontal colored lines show the specific durations for different phases of the rupture sequence, with blue lines indicating slow phases and red lines indicating fast phases. Nucleation and afterslip phases are indicated by two blue lines at the beginning and end of fast events. The colormap represents slip rate. Black circles represent events detected on the fractures, and projected onto the main fault strike. Yellow stars represent the epicenters of the different events. Panel a) shows complex partial ruptures on the main fault. Panel b) shows one slow slip event. Panel c) shows a full rupture on the main fault. Panel d) shows an event, nucleated from one end of the fault and accelerated in a cascading process.
@@
Having explored temporal dynamics, we now examine spatial complexity of the seismic events recorded on the main fault (Figure 3). We delve into the spatial complexity of seismic events recorded on the main fault by looking at the slip rate profiles along strike as a function of spatially averaged cumulative slip during the event. In addition, we project all off-fault activity during this event onto the main fault and denote them by simple black circles. For our analysis, we carefully selected several representative events from the case study, encompassing a diverse range of slip occurrences. This includes partial ruptures, full ruptures, and slow slip events, each characterized by distinct magnitudes and durations. For the fast ruptures, we present both the nucleation phase, where the slip rate starts to accelerate from an arbitrarily small threshold, i.e., $10^{-9}$ m/s, prior to reaching some geodetically detectable threshold, i.e., $10^{-6}$ m/s, and the afterslip phase, where the slip rate decelerates to the same limits. Figure 3a features two partial ruptures, each encompassing around half of the main fault. Notably, the second rupture (moment magnitude $M_{w}=4.70$) nucleates at the edge of the first one ($M_{w}=4.56$), demonstrating a cascading effect. Despite the rapid rupture of coseismic phases, which last 17 and 22 seconds, respectively, the nucleation phase lasting 122 days and the afterslip phase taking place over 30 days are significantly slower.